MegaPrime/MegaCam history in view of data quality

This page aims at synthesizing in a single location all the instrumental evolutions of the MegaPrime/MegaCam setup that had an impact on the data since February 2003 (beginning of semester 03A).

A lot of improvements were made during the commissioning phase between December 2003 and April 2003. However, some light leaks were contaminating the data during that period in the i' and z' bands and also not one exposure was guided. Guiding became operational in May, however, at that time half the mosaic could not be read out for the whole duration of the observing run.

MegaCam was in its full operational mode for the first time in June 2003. Data users wanting to avoid headaches due to a large variety of data of various nature should in consequence focus only on data taken after the 21st of June 2003 (03AQ05 and on). On the other side, data up to the end of the February run should be considered marginal.

Major incidents, or collection of incidents (like in 04BQ01) are reported on the CFHTLS news pages as they relate to data quality affecting the survey data in particular, but the notes apply to all data.

All times in these tables are in Hawaiian local time. A start date defines the beginning of a night and an end date defines the end of night (the problem usually getting fixed during the day), e.g. "05 Sep. 04 - 07 Sep. 04" means from the evening of the 5th of Sep. to the morning of the 7th.

All effects since June 2003 (03AQ05 and on only)
Time rangeQSO run(s)EventEffect on data
30 Mar. 16 - 14 Apr. 16 16AQ07 (first 2 nights, last night) Chip #25 amp A gain variation Photometry is affected in that amp.
03 Dec. 15 - 07 Dec. 15 noon 15BQ14 (first part) Faulty power supply (slave board) Unusable lower half of the mosaic
02 Oct. 15 - present 15BQ08 and on Astrometry based on scamp software, linear distorsion model Avoid chips with no astrometric solution
04 Sep. 15 - present 15BQ06 and on First zero points for narrow, ultra-broad band filters Based on HST spectrophotometric standards
09 Jun. 15 - present 15AQ10 and on Changed from synthetic to empirical color corrections New broad band filters only
10 Apr. 15 - present 15AQ06 and on Changed mode to median in master flat creation Fixed master flats for bimodal distributions
10 Apr. 15 - present 15AQ06 and on New photometric grid solutions Valid for new filters, on all 40 CCDs
10 Feb. 15 - present 15AQ02 and on New photometric calibration Zero point estimates based on SNLS catalogs
10 Feb. 15 - present 15AQ02 and on New broad and narrow band filters. Ear CCDs are not vignetted
10 Feb. 15 - present 15AQ02 and on 40 CCDs images. Images now have 40 extensions.
12 Dec. 14 - 30 Dec. 14 14BQ11 Degraded L1 coating Very large scattering halos, 50% flux lost in PSF core
25 Sep. 13 - 08 Jan. 14 13BQ05 to 13BQ11 Slow drift in zero points (up to 0.2 mag in u)Only throughput was impacted
18 Oct. 07 - present 07BQ06 and onNew i' filter installed (i2)Slightly broader and bluer filter
20 Jul. 07 - n/a 05BQ10 Original i' filter broken Filter can not be perfectly duplicated
19 May - 27 Dec.06 06AQ08-06BQ08 Slow growth of metal shading Attenuation gradient (5 to 15%)
20 Sep. 05 - present 05BQ05 and on Tweak CCD focal plane tilt More uniform IQ across focus
01 Sep. - 02 Sep.05 05BQ03 Slave power supply failure Lower half mosaic down (1 night)
20 Jul. 05 - present 05AQ07 and on 5.5 mm -> 2.1 mm spacer More uniform IQ, less distortion
07 Jul. 05 - present 05AQ06 and on Tighter material near mirror Eliminate 0.5% flux variation vs.dec
02 Dec. 04 - present 04BQ05 and on L3 lens flipped upside down Spectacular IQ improvement
01 Nov.- 04 Nov.04 04BQ04 L3 lens flipped upside down Spectacular IQ improvement
08 Sep.- 22 Oct.04 04BQ02-Q03 Defective video board MAXLINB=32K right amp CCD27->35
05 Sep.- 07 Sep.04 04BQ02 Defective video board Half mosaic not operational (18-35)
14 Aug.- 24 Aug.04 04BQ01 CCDs 00,11 defective Only A amplifier (left) off for 07&11
06 Aug.- 14 Aug.04 04BQ01 CCDs 00,07,09,10,11 defective Only A amplifier (left) off for 00&11
16 Aug.- 17 Aug.04 04BQ01 Ice on window dusk to 1 AM Large blob obstructing light
06 Aug.- 07 Aug.04 04BQ01 Ice on window all night Large blob obstructing light
06 Aug.- 10 Aug.04 04BQ01 Vacuum gauge glow Large structures in all bands
07 Jul. 04 - present 04AQ07 and on L1 cell O-ring removal No obvious IQ improvement
25 Feb. 04 - present 04AQ01 and on Light baffle installed Cut stray lights
13 Jan.- 30 Jan.03 03BQ07 CCD03 non operational 2.7% off the field of view at the edge
16 Dec.- 31 Dec.03 03BQ06 CCD03 non operational 2.7% off the field of view at the edge
28 Nov.- 03 Dec.03 03BQ05 CCD03 non operational 2.7% off the field of view at the edge
27 Oct.- 01 Nov.03 03BQ04 CCD03 non operational 2.7% off the field of view at the edge
22 Oct.- 22 Oct.03 03BQ04 Ice on window 3 AM to dawn Large blob obstructing light
18 Sep.- 04 Oct.03 03BQ03 CCD03 non operational 2.7% off the field of view at the edge
18 Jul. 03 - Jul.05 03AQ06-present WFC original tilt & centering
with 5.5mm spacer
Back to best optical setup
16 Jun.- 17 Jul.03 03AQ05 WFC 0.02 deg. tilt to west
side with 5.5mm spacer
Not optimal optical setup yet
Camera
Time rangeQSO run(s)EventEffect on data
30 Mar. 16 - 14 Apr. 16 16AQ07 (first 2 nights, last night) Chip #25 amp A gain variation Photometry is affected in that amp.
03 Dec. 15 - 07 Dec. 15 noon 15BQ14 (first part) Faulty power supply (slave board) Unusable lower half of the mosaic
10 Feb. 15 - present 15AQ02 and on New broad and narrow band filters. Ear CCDs are not vignetted
10 Feb. 15 - present 15AQ02 and on 40 CCDs images. Images now have 40 extensions.
18 Oct. 07 - present 07BQ06 New i' filter installed (i2)Slightly broader and bluer filter
20 Jul. 07 - n/a 05BQ10 Original i' filter broken Filter can not be perfectly duplicated
01 Sep. 05 - 02 Sep. 05 05BQ03 Power supply failure Lower half of the mosaic down (1 night)
08 Sep. 04 - 22 Oct. 04 04BQ02&03 Defective video board MAXLINB=32K right amp CCD27->35
05 Sep. 04 - 07 Sep. 04 04BQ02 Defective video board Half mosaic not operational (18-35)
14 Aug. 04 - 24 Aug. 04 04BQ01 CCDs 00,11 defective Only A amplifier (left) off for 07&11
06 Aug. 04 - 14 Aug. 04 04BQ01 CCDs 00,07,09,10,11 defectiveOnly A amplifier (left) off for 00&11
13 Jan. 03 - 30 Jan. 03 03BQ07 CCD03 non operational 2.7% off the field of view at the edge
16 Dec. 03 - 31 Dec. 03 03BQ06 CCD03 non operational 2.7% off the field of view at the edge
28 Nov. 03 - 03 Dec. 03 03BQ05 CCD03 non operational 2.7% off the field of view at the edge
27 Oct. 03 - 01 Nov. 03 03BQ04 CCD03 non operational 2.7% off the field of view at the edge
18 Sep. 03 - 04 Oct. 03 03BQ03 CCD03 non operational 2.7% off the field of view at the edge
21 May 03 - 29 May 03 03AQ04 Power supply failure Low 50% CCD mosaic not operational
Wide-field corrector (in 2003)
Time rangeQSO run(s)EventEffect on data
12 Dec. 14 - 30 Dec. 14 14BQ11 Degraded L1 coating Very large scattering halos, 50% flux lost in PSF core
20 Sep.- present 05BQ05-presentTweak MegaCam focal plane tilt More uniform IQ across focus
20 Jul.- present 05AQ07-present5.5 mm -> 2.1 mm spacer More uniform IQ, less distortion
02 Dec. 04 - present 04BQ05-presentL3 lens flipped upside down Spectacular IQ improvement
01 Nov. - 04 Nov. 04 04AQ04 WFC's L3 lens flipped upside down Spectacular IQ improvement
07 Jul. 04 - present 04AQ07-presentL1 cell O-ring removal No obvious improvement
18 Jul. 03 - present 03AQ06-presentOriginal tilt & centering Back to best setup
16 Jun.- 17 Jul. 03 03AQ05 0.02 deg. tilt to west side A bit worse actually
15 May - 15 Jun. 03 03AQ04 Original tilt & centering Best configuration
28 Apr.- 14 Jun. 03 03AQ03 WFC moved west by 1 mm Worse
16 Apr. 03 - Jul. 06 03AQ03 7 mm -> 5.5mm spacer, 0.05 deg. tilt Getting better
25 Jan.- 15 Apr. 03 03AQ01-03AQ02 7 mm spacer + focal plane tilts Significant improvement
12 Dec.- 24 Jan. 03 n.a. Original configuration Marginal image quality
Baffling
Time rangeEventEffect on data
Jul. 05Tightened blackening materialEliminate a possible 0.5% flux variation depending of DEC
Feb. 04Light baffle installedNo more stray lights, reduced sky backgrounds
Aug. 03Sides of mirror covers painted blackNo more stray lights from the Moon
Feb. 03Telescope white surfaces blackenedSerious decrease in large gradients when the Moon is up
Guiding
Time rangeQSO run(s)EventEffect on data
Jun. 03 - present 03AQ05 - presentFine tune-upGuiding/tip-tilt optimized
May 03 - Jun. 0303AQ04Initial releaseGuiding not yet optimized
Dec. 02 - Apr. 0303AQ01 - 03AQ03No guidingOnly short tracked exposures taken
Primary mirror wash & recoating
Time rangeQSO run(s)EventMeasured reflectivity
08/24/07 - present 07BQ04 - presentRecoatingTBD
12/15/05 - 08/21/07 05BQ11 - 07BQ02 Wash88.2%
08/31/05 - 12/15/05 05BQ03 - 05BQ09 WashTBD
01/26/05 - 08/31/05 05AQ01 - 05BQ03 Wash89.72%
04/14/04 - 01/26/05 04AQ03 - 04BQ06 Wash88.79%
08/15/03 - 04/14/04 03AQ06 - 04AQ02 Recoating87.98%
04/17/02 - 08/15/03 First light - 03AQ05 Wash85.50%.
Internal light leaks
Time rangeQSO run(s)EventEffect on data
May 19 to Sep. 1 200606AQ08-06BQ01Infrared LED leakVariable structure only in z'
Aug. 6 to 10 (AM) 200404BQ01Vacuum gauge glowLarge structures in all bands
Dec. 2002 to Apr. 200303AQ01-03AQ03Infrared LED leakLarge structures in i' and z'
Ice & dust on cryostat window
Time rangeQSO run(s)EventEffect on data
19 May - 27 Dec. 06 06AQ08-06BQ08 Slow growth of metal shading Attenuation gradient (5 to 15%)
16 Aug. - 17 Aug. 04 04BQ01 Ice on window dusk to 1 AM Large blob obstructing light
06 Aug. - 07 Aug. 04 04BQ01 Ice on window all night Large blob obstructing light
22 Oct. - 22 Oct. 03 03BQ04 Ice on window 3 AM to dawn Large blob obstructing light